A small turbine using compressed air has a power output of 100 W.




The enthalpy of the air at the inlet is 320 kJ/kg and the enthalpy at the outlet is 240 kJ/kg. Assume the process is adiabatic and the change in the kinetic energy is negligible. Compute the mass flow rate.

1
Expert's answer
2022-04-22T13:53:38-0400

"\dot m=\frac{S_1-S_2}W=0.8~\frac{kg}s ."